OpenStudy (anonymous):

Find the horizontal or oblique asymptote of f(x) = 2x^2 + 5x + 6/x + 3

OpenStudy (anonymous):

Answer choices:

OpenStudy (anonymous):

@Destinymasha @phi @Punkchick @Luigi0210 @Callisto @marylou004 @emily_grace2 Help me out please :)

OpenStudy (anonymous):

do long devision

OpenStudy (anonymous):

How? lol

OpenStudy (anonymous):

@satellite73 @ehuman @wolf1728 Can you help me out please?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

That's all that it asks.

OpenStudy (anonymous):

Okay, thanks :)

OpenStudy (anonymous):

Perhaps @broken_symmetry @superdavesuper @dresdendrew can help :)

OpenStudy (anonymous):

Okay, so how would I get the answer after that?

OpenStudy (anonymous):

@satellite73 Help?

OpenStudy (anonymous):

For the function you quoted, there is one vertical asymptote at x=0 and no others. Did you mean \[y = \frac{2x^2 + 5x + 6}{x + 3}\] instead of \[y = 2x^2 + 5x + \frac6x + 3\]?

OpenStudy (anonymous):

Yes, I did

OpenStudy (anonymous):

@broken_symmetry

OpenStudy (anonymous):

Have you tried dividing the polynomial? Try to get something of the form \[(Ax + B)(x+3) = 2x^2 + 5x+6-C\] This will mean that \[y = (Ax+B)+ \frac{C}{x+3}\] The function Ax+B will be the asymptote you're looking for

OpenStudy (anonymous):

Okay, I'm following. Can you show me how to get the answer? :)

OpenStudy (anonymous):

Multiply out (Ax+B)(x+3) Then compare coefficients of x^2, x and 1 with the right hand side

OpenStudy (anonymous):

So, it'll be y = -5? I haven't done this stuff for weeks, I'm probably wrong :/

OpenStudy (anonymous):

No, what I mean is do \[(Ax+B)(x+3) = Ax^2+(B+3A)x+3B = 2x^2 + 5x + 6 - C\] This tells you A = 2 3A+B = 5 3B = 6-C The function Ax+B is your answer because if \[y = (Ax+B)+\frac{C}{x+3}\] then \[\lim_{x\to\pm\infty} y \to Ax+B\]

OpenStudy (anonymous):

Oh okay, so did I get it right? What's the answer? :)

OpenStudy (anonymous):

Can you please just give me the answer? I've been working on this one problem for an hour!

OpenStudy (luigi0210):

Figure it out?